# Holographic analysis of near-conformal dynamics and light dilaton

### v1

@@ 4.2.2 The flavor field and the Efimov spiral
We start by discussing the behavior near the boundary, $r\sim r_{\mathrm{UV}}$. The first observation about the background is that the equation of motion for $X$ is linear. In our case this appears to be the case by default, since the action (73) is quadratic in $X$, but this is expected to hold in general: $X$ is small for $r\ll r_{\mathrm{IR}}$, and the nonlinear terms are negligible. Actually, we will find $X(r\sim r_{\mathrm{UV}})\propto(r_{\mathrm{UV}}/r_{\mathrm{IR}})^{2}$ which reflects the IR dimension $\Delta_{\mathrm{IR}}$ for $\nu\to 0$. The nonlinear terms will only become important in the IR, $r\sim r_{\mathrm{IR}}$, and in part determine the IR dynamics which we parametrize in this work through boundary conditions. Moreover notice that even though the action (73) is linear in $X$, nonlinearities could still arise due to the backreaction to the geometry. Therefore it is important even in the current case that $X$ is suppressed in the region that we are analysing.
Due to linearity and the smallness of $X$, the solution to (75) can be written as the sum of the source and vev terms: [^4] For generic actions the source term is only well defined up to a multiple of the vev term, see, e.g. Jarvinen:2015ofa . This ambiguity will not affect the analysis in this article.
@@
where $i=1,2$. That is, the flow of the solution for $r\sim r_{\mathrm{UV}}$ implies a nontrivial transformation between the bases that are natural near the boundary ($r\ll r_{\mathrm{UV}}$) and the walking regime $r_{\mathrm{UV}}\ll r\ll r_{\mathrm{IR}}$. The coefficients $C_{i}^{(X)}\cos\rho_{i}$ and $C_{i}^{(X)}\sin\rho_{i}$ form a transformation matrix between the bases which is nontrivial due to the flow. Note that both (86) and (4.2.2) are valid in the region ($r\ll r_{\mathrm{IR}}$) where $X$ is small and nonlinear terms in its equation of motion can be safely neglected, as we have done here. This guarantees that the transformation is linear. Moreover, the definition of the coefficients $C_{i}^{(X)}$ and $\rho_{i}$ makes no reference to the values of $m_{q}$ and $\sigma$, so the coefficients are independent of these parameters.
@@ 4.3 Fluctuations and light modes
where $g_{\mu\nu}$ are the space-time components of the metric, and $\varphi$, $\chi$ and $\psi$ are the infinitesimal fluctuations. As we will take the fluctuations to be plane waves in space-time coordinates with nonzero mass, the inverse of $\partial^{2}$ is well-defined here. To proceed, we can define the following diffeormorphism invariant combinations:
@@ 4.3.1 Two-point correlation function of the quark bilinear and the UV completion
Let us then analyze the two-point functions. We first focus on the two-point function of the quark bilinear $\bar{\psi}\psi$, which is the easiest correlator to analyze. Also, it is expected to show the same potentially light mode as the correlator of the trace of the energy-momentum tensor $\theta_{\mu}^{\mu}$. We also discuss how the result for the two-point functions can be expressed in terms of coefficients of the solutions in the walking regime. This allows us to interpret how the results of the UV-completed setup are linked to simpler models where the UV completion is neglected.
@@
In the notation of Section 4.2, the exact source and vev include additional nontrivial phase factors $\rho_{i}$ (see, e.g., (4.2.2)). The phase factors were defined for the background $X$, but as we argued above, the field $\hat{\xi}(r)$ satisfies the same equation as the background for $r\ll r_{\mathrm{IR}}$ so long as the frequency of the fluctuation is small, $\omega\sim 1/r_{\mathrm{IR}}$. Therefore, the (similarly defined) phase factors for the fluctuations exactly match with those of the background. To use this information, we may first rewrite (115) as
@@
which holds for $r_{\mathrm{UV}}\ll r\ll r_{\mathrm{IR}}$ if $\omega\sim 1/r_{\mathrm{IR}}$. Here the first (second) term gives the source (vev) of the UV-completed setup. Therefore the exact correlator of the bilinear $\bar{\psi}\psi$ reads in terms of the data $\Xi_{\pm}$ from the walking region
@@
We then compare the exact result to some simple approximations which do not use the any information about the UV-completion but only the solutions in the walkig region. The oscillating solution of (115) may also be written as
@@
However, since it is expected that the system does will have a separate UV scale, it may be obvious that source and vev definitions based solely on IR behavior are too simplistic. Indeed, the approximation (123) does not seem to reflect the phase structure of the exact result (120) – recall that the natural expectation for the phases $\rho_{i}$, which is also supported by the model of Sec. 3, is that they are small, $\rho_{i}\sim\nu$.
@@
Notice that requiring the vev term to vanish at the cutoff does not uniquely determine the source function in (4.3.1). It does however determine the coefficient of the source up to a trivial constant. This is the most important piece in the correlator (125) as it appears in the denominator and controls the poles of the correlator, i.e., the spectrum. Redefinitions of the source function only modify the correlator such that the pole structure is not altered.
@@ 4.3.3 Correlator of the trace of the energy momentum tensor
Let us then comment on the correlator of $\theta_{\mu}^{\mu}$. Similarly as the one-point function considered in Appendix E.1, the correlator has contributions from both from the gluon and from the flavor sector. The contribution from the flavor sector is expected to have similar structure as the correlator of the quark bilinear discussed above. In particular, it contains poles due to the light mode if such a mode is present. However, it is tricky to write compact expressions for the general correlator such as (120) in this case. The main reason is that, computing the correlator of $\theta_{\mu}^{\mu}$ requires adding sources both for the gauge invariant fields $\xi$ and $\zeta$ in (108) and in (109) (see Appendices B and E.1). The mass term in the $\xi$-equation (108) cannot be neglected, which leads to a nontrivial coupling between the fields. Therefore, we do not try to analyze this correlator precisely in the UV-completed setup of this section.
However, we can make interesting remarks. Namely, in the limit $\omega\to 0$, taking $\xi=\zeta=\mathrm{const.}$ gives an exact solution to (108) and (109) [^6] Note that strictly speaking, the decomposition of the gravity fluctuation modes in Appendix B only makes sense at nonzero mass of the fluctuations. The case of zero mass is special Kiritsis:2006ua . Therefore this massless solution only makes sense as the limiting solution $\omega\to 0$ of the massive solution.. While we do not specify the action in the IR region, we expect that reasonable choices lead to this being the precise IR-normalizable solution. This means that the solution only has source terms in the limit $\omega\to 0$, so that the correlator must vanish in this limit. Further, the IR boundary condition in our language (e.g., that given in (126)), which picks this solution is the Neumann condition, $\mathcal{A}=0$. That is, apart from leading to light modes, the Neumann boundary condition is singled out by this particular constant solution. Since the Neumann boundary conditions in general seem to lead to light modes, this gives further support to the presence of such modes in a properly IR completed setup.
@@ 4.4 Flavor effects in Green’s functions and identities
where we used the identity in (100) to obtain the latter expression. That is, this identity shows that the latter term is negative as was the case for the toy model in (56). Therefore at zero quark mass, the solution with the smallest amount of scale separation (i.e., largest $r_{\mathrm{IR}}$) has always the lowest free energy, and as in the case of the toy model the same is typically found also for nonzero mass. In (135) we also included the constant of integration $F_{0}$, which is interpreted as the free energy of the solution in the origin of the spiral. This solution describes a flow from the UV fixed point to the IR fixed point, which is unperturbed by the flavors. Moreover, it may appear surprising that the result for the free energy depends on the IR scale $r_{\mathrm{IR}}$, which was defined in a somewhat arbitrary way in (84). Note however that the dependence is only through the combination $X_{\mathrm{IR}}/r_{\mathrm{IR}}^{2}$, which is actually the combination of (90) where the dependence on the definitions of $r_{\mathrm{IR}}$ cancels.
@@
$\langle\bar{\psi}\psi\rangle=\frac{\partial F}{\partial m_{q}}=-\frac{2}{\kappa^{2}}\mathcal{N}\ell^{3}\sigma\ ,$ (138)
where the last expression could be read off from (134) or computed using the dictionary following the steps in Appendix E. In Appendix F we also analyze the latter term of (136), finding that
$-\frac{\ell^{3}}{2\kappa^{2}}(4-\Delta_{g})\Delta_{g}\phi_{s}\phi_{v}=\frac{N_{c}}{8\lambda^{2}}\beta(\lambda)\langle\mathbb{T}r\,G_{\mu\nu}G^{\mu\nu}\rangle\ ,$ (139)
@@
$\langle\theta^{\mu}_{\mu}\rangle=m_{q}\langle\bar{\psi}\psi\rangle+\frac{N_{c}}{2\lambda^{2}}\beta(\lambda)\langle\mathbb{T}r\,G_{\mu\nu}G^{\mu\nu}\rangle\ .$ (140)
@@ 5 Conclusion
We have studied the dynamics of gauge theories near the conformal edge in the context of gauge-gravity duality. In particular, we have analyzed in detail the properties of the PNGB dilaton, which is generated when the scale symmetry of the near-conformal gauge theory is spontaneously broken by condensations of a quark-bilinear and gluons as well. For our analysis, we considered a gravity dual based on a five-dimensional geometry, which is asymptotically Anti-de Sitter in the deep UV. We also included bi-fundamental bulk scalars and a dilaton field, dual to the quark bilinears, and the trace of the gluon field strength squared, respectively.
When the gauge theory is near conformal, the anomalous dimension of the quark bilinear is close to one. This is reflected in the analysis of the gravity dual such that the bulk scalar violates the BF bound slightly. The IR geometry of the gravity dual for confining gauge theories, in general, deviates a lot from the AdS space, capping off in the IR for example. We cut off the space at the IR boundary where the walking behavior stops, while introducing a general, but linear, boundary condition at the IR boundary for the bulk scalars. We found that the PNGB dilaton is parametrically light compared to all other hadrons when the IR boundary condition is almost Neumann. The PNGB dilaton is a diffeomorphism-invariant mixture of a bulk scalar and the dilaton (or the scalar graviton). However, we find that the light mode is dominantly driven by the fluctuations of the scalar, suggesting that it is mostly a mesonic (rather than a glueball) state. This is in agreement with previous studies in related models Kutasov:2011fr ; Arean:2013tja ; Pomarol:2019aae . We have checked for consistency that the PNGB dilaton saturates the anomalous Ward identity of the scale symmetry. We also found that the hypothesis of the partially conserved dilatation current (PCDC) imposes a rather strong constraint on IR parameters to be consistent with the scale anomaly calculation in field theory by Gusynin and Miransky. We also calculated, for the case of a small quark mass, the pion mass and its decay constant in the gravity dual to find that the GOR relation holds.
@@
However, IR physics is, in general, more complicated. In the holographic model, the scalar fields grow large in the IR region, meaning that nonlinearities and backreaction effects become important. This reflects the complexity of the IR physics in field theory. In this work, we implemented the IR physics by considering generic IR boundary conditions for the fields and their fluctuations. We argued that such boundary conditions are indeed capable of capturing generic IR physics. We observed, in agreement with earlier literature (see, e.g., Arean:2013tja ; Elander:2018gte ), that while large scale separation leads to special scalar modes, whether they are (anomalously) light in the end depends on the IR boundary conditions.
It would be important to check that our findings hold in a model that (apart from the UV flow) explicitly implements an IR completion following some of the examples in the literature, for example Jarvinen:2011qe ; Kutasov:2012uq ; Pomarol:2019aae . Such an IR completion should be general enough to capture all possibilities considered here, in particular, the cases which show large scale separations between the UV and IR that differ from the standard “Miransky” scaling law. Note that it was in such cases with nonstandard scaling, and with Neumann boundary conditions for the fluctuations, where we found an anomalously light scalar mode in our analysis. Explicit modeling of the IR physics should confirm that it is indeed possible to obtain this kind of mixture (both for the background and the fluctuations). We already presented an argument based on the analysis of the fluctuation equations in Sec. 4.3 that the Neumann boundary conditions will be singled out by a full model, but this should also be confirmed by explicit solutions. Having a full model of the IR dynamics would also allow us to study the mixing of the meson and glueball modes in the IR region explicitly.
@@ Appendix A Precise dictionary between gravity and field theory
In this Appendix, we review the dictionary between the holographic model and the gauge theory. We first construct the relevant boundary term on the QCD side of the gauge/gravity correspondence, which fixes how the source terms of the gravity couple to the operators of field theory. We start by considering the QCD action generalized to a background with the metric $\gamma$ which is assumed to be close to the Minkowski metric:
$S_{\gamma}=\int d^{4}x\,\sqrt{-\det\gamma}\bigg{[}-{1\over 2g^{2}}\,{\mathbb{T}r}G_{\mu\nu}G_{\rho\sigma}\gamma^{\mu\rho}\gamma^{\nu\sigma}+i\gamma^{\mu\nu}\bar{\psi}\gamma_{\mu}D_{\nu}\psi-\bar{\psi}_{R}\,M_{q}\,\psi_{L}-\bar{\psi}_{L}\,M_{q}^{\dagger}\,\psi_{R}\bigg{]}\,,$ (146)
where $\psi_{L}=(1+\gamma^{5})\,\psi/2$, $\psi_{R}=(1-\gamma^{5})\,\psi/2$, and $M_{q}$ is the (potentially complex) quark mass matrix.
@@
We also need the dictionary for the other operators. These are the dilaton field $\phi$ dual to the ${\mathbb{T}r}\,G_{\mu\nu}^{2}$, and the complex scalar $X$ which is a matrix in flavor space and dual to $\bar{\psi}\psi$. The simplest possibility is that for asymptotically AdS5 spaces the dilaton approaches a constant at the boundary, which is identified as the source for the ${\mathbb{T}r}\,G_{\mu\nu}^{2}$ operator, i.e., the gauge coupling. If the dimension of the operator $\Delta_{g}$ is less than four, we can write $\phi(z,x)=\phi_{0}(x,z)+{\cal O}\left(z^{2\Delta_{g}-4}\right)=\phi_{s}(x)z^{4-\Delta_{g}}+{\cal O}\left(z^{2}\right)$. We consider more complicated cases below in Appendix F. The flavor field $X$ is assumed to have the asymptotics near the boundary
@@
$S_{\mathrm{bdry}}\!\!$ $=$ $\!\!\int d^{4}x\bigg{[}-\frac{1}{2}\theta^{\mu\nu}\delta\hat{g}_{\mu\nu}-\frac{N_{c}}{2}e^{-\phi}\ {\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}-\frac{K_{X}}{z^{\Delta-4}}\bar{\psi}_{R}\,X\,\psi_{L}-\frac{\bar{K}_{X}}{z^{\Delta-4}}\bar{\psi}_{L}\,X^{\dagger}\,\psi_{R}\bigg{]}_{z=\epsilon}$ (154)
$=$ $\!\!\int d^{4}x\bigg{[}-\frac{1}{2}\theta^{\mu\nu}\delta\gamma_{\mu\nu}-\frac{N_{c}}{2}e^{-\phi_{0}}\ {\mathbb{T}r}G_{\mu\nu}G^{\mu\nu}-K_{X}\bar{\psi}_{R}\,X_{-}\,\psi_{L}-\bar{K}_{X}\bar{\psi}_{L}\,X_{-}^{\dagger}\,\psi_{R}\bigg{]}$
@@
The interpretation is that the UV cutoff $\epsilon$ maps to the (inverse of the) renormalization scale of field theory. Here $\delta\hat{g}_{\mu\nu}=\hat{g}_{\mu\nu}-\eta_{\mu\nu}$ and $K_{X}$ is a constant whose value cannot be determined. This reflects the fact that the quark mass is not an observable. In the analysis in this article we will set the constant to be one for simplicity. Note that the latter three terms also formally contain the kinetic term of the gauge field as well as the fermion mass terms in the QCD action. In particular, we identify $X_{-}$ as the quark mass matrix, $M_{q}=X_{-}$. Below we will only consider the case where both $X_{-}$ and $M_{q}$ are proportional to the identity matrix in flavor space. In this case we denote the (real) quark mass as $m_{q}$.
Let us then comment on the five dimensional gravity action. It is given by various term as follows:

### v2

@@ 4.2.2 The flavor field and the Efimov spiral
We start by discussing the behavior near the boundary, $r\sim r_{\mathrm{UV}}$. The first observation about the background is that the equation of motion for $X$ is linear. In our case this appears to be the case by default, since the action (75) is quadratic in $X$, but this is expected to hold in general: $X$ is small for $r\ll r_{\mathrm{IR}}$, and the non-linear terms are negligible. Actually, we will find $X(r\sim r_{\mathrm{UV}})\propto(r_{\mathrm{UV}}/r_{\mathrm{IR}})^{2}$ which reflects the IR dimension $\Delta_{\mathrm{IR}}$ for $\nu\to 0$. The non-linear terms will only become important in the IR, $r\sim r_{\mathrm{IR}}$, and in part determine the IR dynamics which we parametrize in this work through boundary conditions. Moreover notice that even though the action (75) is linear in $X$, non-linearities could still arise due to the backreaction to the geometry. Therefore it is important even in the current case that $X$ is suppressed in the region that we are analysing.
Due to linearity and the smallness of $X$, the solution to (77) can be written as the sum of the source and vev terms: [^6] For generic actions the source term is only well-defined up to a multiple of the vev term, see, e.g. Jarvinen:2015ofa . This ambiguity will not affect the analysis in this article.
@@
where $i=1,2$. That is, the flow of the solution for $r\sim r_{\mathrm{UV}}$ implies a non-trivial transformation between the bases that are natural near the boundary ($r\ll r_{\mathrm{UV}}$) and the walking regime $r_{\mathrm{UV}}\ll r\ll r_{\mathrm{IR}}$. The coefficients $C_{i}^{(X)}\cos\rho_{i}$ and $C_{i}^{(X)}\sin\rho_{i}$ form a transformation matrix between the bases which is non-trivial due to the flow. Note that both (88) and (4.2.2) are valid in the region ($r\ll r_{\mathrm{IR}}$) where $X$ is small and non-linear terms in its equation of motion can be safely neglected, as we have done here. This guarantees that the transformation is linear. Moreover, the definition of the coefficients $C_{i}^{(X)}$ and $\rho_{i}$ makes no reference to the values of $m_{q}$ and $\sigma$, so the coefficients are independent of these parameters.
@@ 4.3 Fluctuations and light modes
where $g_{\mu\nu}$ are the space-time components of the metric, and $\varphi$, $\chi$ and $\psi$ are the infinitesimal fluctuations. As we will take the fluctuations to be plane waves in space-time coordinates with non-zero mass, the inverse of $\partial^{2}$ is well-defined here. To proceed, we can define the following diffeormorphism invariant combinations:
@@ 4.3.1 Two-point correlation function of the quark bilinear and the UV completion
Let us then analyze the two-point functions. We first focus on the two-point function of the quark bilinear $\bar{q}q$, which is the easiest correlator to analyze. Also, it is expected to show the same potentially light mode as the correlator of the trace of the energy-momentum tensor $\theta_{\mu}^{\mu}$. We also discuss how the result for the two-point functions can be expressed in terms of coefficients of the solutions in the walking regime. This allows us to interpret how the results of the UV-completed setup are linked to simpler models where the UV completion is neglected.
@@
In the notation of Section 4.2, the exact source and vev include additional non-trivial phase factors $\rho_{i}$ (see, e.g., (4.2.2)). The phase factors were defined for the background $X$, but as we argued above, the field $\hat{\xi}(r)$ satisfies the same equation as the background for $r\ll r_{\mathrm{IR}}$ so long as the frequency of the fluctuation is small, $\omega\sim 1/r_{\mathrm{IR}}$. Therefore, the (similarly defined) phase factors for the fluctuations exactly match with those of the background. To use this information, we may first rewrite (117) as
@@
which holds for $r_{\mathrm{UV}}\ll r\ll r_{\mathrm{IR}}$ if $\omega\sim 1/r_{\mathrm{IR}}$. Here the first (second) term gives the source (vev) of the UV-completed setup. Therefore the exact correlator of the bilinear $\bar{q}q$ reads in terms of the data $\Xi_{\pm}$ from the walking region
@@
We then compare the exact result to some simple approximations which do not use any information about the UV-completion but only the solutions in the walkig region. The oscillating solution of (117) may also be written as
@@
However, since it is expected that the system will have a separate UV scale, it may be obvious that source and vev definitions based solely on IR behavior are too simplistic. Indeed, the approximation (125) does not seem to reflect the phase structure of the exact result (122) – recall that the natural expectation for the phases $\rho_{i}$, which is also supported by the model of Sec. 3, is that they are small, $\rho_{i}\sim\nu$.
@@
Notice that requiring the vev term to vanish at the cutoff does not uniquely determine the source function in (4.3.1). It does, however, determine the coefficient of the source up to a trivial constant. This is the most important piece in the correlator (127) as it appears in the denominator and controls the poles of the correlator, i.e., the spectrum. Redefinitions of the source function only modify the correlator such that the pole structure is not altered.
@@ 4.3.3 Correlator of the trace of the energy momentum tensor
Let us then comment on the correlator of $\theta_{\mu}^{\mu}$. Similarly as the one-point function considered in Appendix E.1, the correlator has contributions both from the gluon and from the flavor sector. The contribution from the flavor sector is expected to have similar structure as the correlator of the quark bilinear discussed above. In particular, it contains poles due to the light mode if such a mode is present. However, it is tricky to write compact expressions for the general correlator such as (122) in this case. The main reason is that, computing the correlator of $\theta_{\mu}^{\mu}$ requires adding sources both for the gauge invariant fields $\xi$ and $\zeta$ in (110) and in (111) (see Appendices B and E.1). The mass term in the $\xi$-equation (110) cannot be neglected, which leads to a non-trivial coupling between the fields. Therefore, we do not try to analyze this correlator precisely in the UV-completed setup of this section.
However, we can make interesting remarks. Namely, in the limit $\omega\to 0$, taking $\xi=\zeta=\mathrm{const.}$ gives an exact solution to (110) and (111) [^8] Note that strictly speaking, the decomposition of the gravity fluctuation modes in Appendix B only makes sense at non-zero mass of the fluctuations. The case of zero mass is special Kiritsis:2006ua . Therefore this massless solution only makes sense as the limiting solution $\omega\to 0$ of the massive solution.. While we do not specify the action in the IR region, we expect that reasonable choices lead to this being the precise IR-normalizable solution. This means that the solution only has source terms in the limit $\omega\to 0$, so that the correlator must vanish in this limit. Further, the IR boundary condition in our language (e.g., that given in (128)), which picks this solution is the Neumann condition, $\mathcal{A}=0$. That is, apart from leading to light modes, the Neumann boundary condition is singled out by this particular constant solution. Since the Neumann boundary conditions in general seem to lead to light modes, this gives further support to the presence of such modes in a properly IR completed setup.
@@ 4.4 Flavor effects in Green’s functions and identities
where we used the identity in (102) to obtain the latter expression. That is, this identity shows that the latter term is negative as was the case for the toy model in (58). Therefore at zero quark mass, the solution with the smallest amount of scale separation (i.e., largest $r_{\mathrm{IR}}$) has always the lowest free energy, and as in the case of the toy model the same is typically found also for non-zero mass. In (137) we also included the constant of integration $F_{0}$, which is interpreted as the free energy of the solution in the origin of the spiral. This solution describes a flow from the UV fixed point to the IR fixed point, which is unperturbed by the flavors. Moreover, it may appear surprising that the result for the free energy depends on the IR scale $r_{\mathrm{IR}}$, which was defined in a somewhat arbitrary way in (86). Note however that the dependence is only through the combination $X_{\mathrm{IR}}/r_{\mathrm{IR}}^{2}$, which is actually the combination of (92) where the dependence on the definitions of $r_{\mathrm{IR}}$ cancels.
@@
$\langle\bar{q}q\rangle=\frac{\partial F}{\partial m_{q}}=-\frac{2}{\kappa^{2}}\mathcal{N}\ell^{3}\sigma\ ,$ (140)
where the summation over the flavor indices of the quark fields $\bar{q}$ and $q$ is assumed, and the last expression could be read off from (136) or computed using the dictionary following the steps in Appendix E. In Appendix F we also analyze the latter term of (138), finding that
$-\frac{\ell^{3}}{2\kappa^{2}}(4-\Delta_{g})\Delta_{g}\phi_{s}\phi_{v}=\frac{N_{c}}{8\lambda^{2}}\beta(\lambda)\langle\mathrm{Tr}\,G_{\mu\nu}G^{\mu\nu}\rangle\ ,$ (141)
@@
$\langle\theta^{\mu}_{\mu}\rangle=m_{q}\langle\bar{q}q\rangle+\frac{N_{c}}{2\lambda^{2}}\beta(\lambda)\langle\mathrm{Tr}\,G_{\mu\nu}G^{\mu\nu}\rangle\ .$ (142)
@@ 5 Conclusion
We have studied the dynamics of gauge theories near the conformal edge in the context of gauge/gravity duality. In particular, we have analyzed in detail the properties of the PNGB dilaton, which is generated when the scale symmetry of the near-conformal gauge theory is spontaneously broken by condensations of a quark-bilinear and gluons as well. For our analysis, we considered a gravity dual based on a five-dimensional geometry, which is asymptotically Anti-de Sitter in the deep UV. We also included bi-fundamental bulk scalars and a dilaton field, dual to the quark bilinears, and the trace of the gluon field strength squared, respectively.
When the gauge theory is near conformal, the anomalous dimension of the quark bilinear is close to one. This is reflected in the analysis of the gravity dual such that the bulk scalar violates the BF bound slightly. The IR geometry of the gravity dual for confining gauge theories, in general, deviates a lot from the AdS space, capping off in the IR for example. We cut off the space at the IR boundary where the walking behavior stops, while introducing a general, but linear, boundary condition at the IR boundary for the bulk scalars. We found that the PNGB dilaton is parametrically light compared to all other hadrons when the IR boundary condition is almost Neumann. The PNGB dilaton is a diffeomorphism-invariant mixture of a bulk scalar and the dilaton (or the scalar graviton). However, we find that the light mode is dominantly driven by the fluctuations of the scalar, suggesting that it is mostly a mesonic (rather than a glueball) state. This is in agreement with previous studies in related models Kutasov:2011fr ; Arean:2013tja ; Pomarol:2019aae . We have checked for consistency that the PNGB dilaton saturates the anomalous Ward identity of the scale symmetry. We also found that the hypothesis of the partially conserved dilatation current (PCDC) imposes a rather strong constraint on IR parameters to be consistent with the scale anomaly calculation in field theory by Gusynin and Miransky. We also calculated, for the case of a small quark mass, the pion mass and its decay constant in the gravity dual to find that the Gell-Mann-Oakes-Renner relation holds.
@@
However, IR physics is, in general, more complicated. In the holographic model, the scalar fields grow large in the IR region, meaning that non-linearities and backreaction effects become important. This reflects the complexity of the IR physics in field theory. In this work, we implemented the IR physics by considering generic IR boundary conditions for the fields and their fluctuations. We argued that such boundary conditions are indeed capable of capturing generic IR physics. We observed, in agreement with earlier literature (see, e.g., Arean:2013tja ; Elander:2018gte ), that while large scale separation leads to special scalar modes, whether they are (anomalously) light in the end depends on the IR boundary conditions.
It would be important to check that our findings hold in a model that (apart from the UV flow) explicitly implements an IR completion following some of the examples in the literature, for example Jarvinen:2011qe ; Kutasov:2012uq ; Pomarol:2019aae . Such an IR completion should be general enough to capture all possibilities considered here, in particular, the cases which show large scale separations between the UV and IR that differ from the standard “Miransky” scaling law. Note that it was in such cases with non-standard scaling, and with Neumann boundary conditions for the fluctuations, where we found an anomalously light scalar mode in our analysis. Explicit modeling of the IR physics should confirm that it is indeed possible to obtain this kind of mixture (both for the background and the fluctuations). We already presented an argument based on the analysis of the fluctuation equations in Sec. 4.3 that the Neumann boundary conditions will be singled out by a full model, but this should also be confirmed by explicit solutions. Having a full model of the IR dynamics would also allow us to study the mixing of the meson and glueball modes in the IR region explicitly.
@@ Appendix A Precise dictionary between gravity and field theory
In this Appendix, we review the dictionary between the holographic model and the gauge theory. We first construct the relevant boundary term on the QCD side of the gauge/gravity correspondence, which fixes how the source terms of gravity couple to the operators of field theory. We start by considering the QCD action generalized to a background with the metric $\gamma$ which is assumed to be close to the Minkowski metric:
$S_{\gamma}=\int d^{4}x\,\sqrt{-\det\gamma}\bigg{[}-{1\over 2g^{2}}\,{\mathrm{Tr}}G_{\mu\nu}G_{\rho\sigma}\gamma^{\mu\rho}\gamma^{\nu\sigma}+i\gamma^{\mu\nu}\bar{q}\gamma_{\mu}D_{\nu}q-\bar{q}_{R}\,M_{q}\,q_{L}-\bar{q}_{L}\,M_{q}^{\dagger}\,q_{R}\bigg{]}\,,$ (148)
where $q_{L}=(1+\gamma^{5})\,q/2$, $q_{R}=(1-\gamma^{5})\,q/2$, and $M_{q}$ is the (potentially complex) quark mass matrix.
@@
We also need the dictionary for the other operators. These are the dilaton field $\phi$ dual to the ${\mathrm{Tr}}\,G_{\mu\nu}^{2}$, and the complex scalar $X$ which is a matrix in flavor space and dual to $\bar{q}q$. The simplest possibility is that for asymptotically AdS5 spaces the dilaton approaches a constant at the boundary, which is identified as the source for the $\mathrm{Tr}\,G_{\mu\nu}^{2}$ operator, i.e., the gauge coupling. If the dimension of the operator $\Delta_{g}$ is less than four, we can write $\phi(z,x)=\phi_{0}(x,z)+{\cal O}\left(z^{2\Delta_{g}-4}\right)=\phi_{s}(x)z^{4-\Delta_{g}}+{\cal O}\left(z^{2}\right)$. We consider more complicated cases below in Appendix F. The flavor field $X$ is assumed to have the asymptotics near the boundary
@@
$S_{\mathrm{bdry}}\!\!$ $=$ $\!\!\int d^{4}x\bigg{[}-\frac{1}{2}\theta^{\mu\nu}\delta\hat{g}_{\mu\nu}-\frac{N_{c}}{2}e^{-\phi}\ \mathrm{Tr}G_{\mu\nu}G^{\mu\nu}-\frac{K_{X}}{z^{\Delta-4}}\bar{q}_{R}\,X\,q_{L}-\frac{\bar{K}_{X}}{z^{\Delta-4}}\bar{q}_{L}\,X^{\dagger}\,q_{R}\bigg{]}_{z=\epsilon}$ (156)
$=$ $\!\!\int d^{4}x\bigg{[}-\frac{1}{2}\theta^{\mu\nu}\delta\gamma_{\mu\nu}-\frac{N_{c}}{2}e^{-\phi_{0}}\ \mathrm{Tr}G_{\mu\nu}G^{\mu\nu}-K_{X}\bar{q}_{R}\,X_{-}\,q_{L}-\bar{K}_{X}\bar{q}_{L}\,X_{-}^{\dagger}\,q_{R}\bigg{]}$
@@
The interpretation is that the UV cutoff $\epsilon$ maps to the (inverse of the) renormalization scale of field theory. Here $\delta\hat{g}_{\mu\nu}=\hat{g}_{\mu\nu}-\eta_{\mu\nu}$ and $K_{X}$ is a constant whose value cannot be determined. This reflects the fact that the quark mass is not an observable. In the analysis in this article we will set the constant to be one for simplicity. Note that the latter three terms also formally contain the kinetic term of the gauge field as well as the fermion mass terms in the QCD action. In particular, we identify $X_{-}$ as the quark mass matrix, $M_{q}=X_{-}$. Below we will only consider the case where both $X_{-}$ and $M_{q}$ are proportional to the identity matrix in flavor space. In this case, we denote the (real) quark mass as $m_{q}$.
Let us then comment on the five-dimensional gravity action. It is given by various terms as follows: